User Tools

Site Tools


VLBI teleconference, 7 April 2009

12.00 Perth, 14.00 Eastern, 16.00 New Zealand


Vic:03 9545 8222
NSW:02 9490 8022
ACT:02 6276 6622
Qld:07 3214 2311
WA:08 6436 8844
SA:08 8273 8122
NT:08 8944 8488
Tas:03 6232 5188

Meeting ID: 2394

Chair: Steven Tingay

Attendance: Tingay, Reynolds, Moin, Bignall, C. Hotan, A. Hotan (Curtin); Tzioumis, Phillips (ATNF Marsfield); Lovell (Hobart); Reid (Mt. Pleasant); Stevens (ATNF Narrabri); Rioja, Dodson (UWA)


  • Reports
    • Feb 2009 VLBI session
    • eVLBI demo to Curtin on March 11.
    • e-VLBI/VLBI/Networking & Operations meetings 10-12 March
    • GMRT-Mopra test observation
    • IYA observations
    • EAVN meeting (Phil)
    • KVN (Richard/Maria)
    • IVS and VLBI2010 (Jim)
  • Update on ASKAP VLBI EoI and ICRAR investment plan (Tingay)
    • Issues re EoI meeting on 15 April.
    • H-maser for ASKAP?
  • Data from TIGO, O'Higgins (USB disks)
  • Status of ARCS data transfer tests to iVEC (Phillips)
  • Update on New Zealand antenna (Gulyaev)
  • Update on Curtin cluster at Narrabri (Phillips/Amy)
  • AuScope Update (Lovell)
  • Observations next semester
  • AOB

Actions arising

(new items in blue)


  • One of Curtin Mark5B+ units to be shipped to AUT when they arrive, plus an invoice to AUT. (Steven) - updated 28/01/09 - Shipping cockup by supplier. Delay in delivery of units. 11/2/09 - in progress. DONE
  • Swinburne archive of correlated VLBI data to be moved to ATNF & Curtin at end of all Swinburne correlation (Steven/Claire) - update 28/01/09 - Swinburne correlation complete. Claire transferring data to Curtin. Will tidy up before sending on to ATNF. Claire to advise ATNF on archive size. 11/2/09 - in progress.
  • Options for disks and data transport:
    • Buying excess disks from Swinburne VLBI pool and/or Mk5 disks. (Tasso/Steven) - update 28/01/09 - Tasso and Steven still discussing. Changed to separating disk pools. All SWIN disks to be returned to Swinburne.
    • Initiate discussion on next generation VLBI systems and disks (Mk5b/Mk5c/of-the-self) for the March LBA meeting. (Tasso/Steven) - update 28/01/09 - to be added to agenda for March meeting. UTas to be debriefed and kept in loop with meeting outcomes. DONE
  • Provide additional advice to PIs on scheduling the ATCA to keep it phased up at 8 GHz and above (Tasso) - update 28/01/09 - Richard to have input regards polarisation properties. 11/2/09 - to be tested with observations in next session. Continuing.
  • List on the Wiki available bands/polarisations at each LBA antenna. Also update “standard bands” (from S2) and define “standard” bands for wide-band recording. (Cormac/Tasso). 11/2/09 - initial progress within SCHED. Will be further followed up.
  • Curtin Cluster at Narrabri: we would like to get the cluster up once the network is installed. Action on Tasso and Shaun to get the networking going as soon as possible (Tasso/Shaun) - update 28/01/09 - Gear ordered and arrived. Work to be done in next ~3 weeks. Jamie at Narrabri maybe get involved.


  • Observing support schedule for next session on the Wiki
  • Expedite ARCS support for data transfers (Steven/Cormac/Tasso/Chris)


  • Next telecon: Thu 21 May, 2009, 2pm Sydney time (Tasso/Chris)
lbaops/telecon2009apr7.txt · Last modified: 2015/12/18 16:39 by